# Movable Antenna Empowered Near-Field Sensing via Antenna Position Optimization

### v1

@@ II System Model
During the sensing process, the MA array transmits sensing signals and receives the echoes reflected from the target, which is assumed to remain static throughout the process[wang2023near, liu2022cramer]. To characterize the near-field channel from the antenna array to the target, we adopt the uniform spherical wave (USW) channel model in [liu2023near], where the channel coefficients have identical amplitudes across all MAs, while their phases vary among them. As depicted in Fig. 1, we denote the physical steering angle from the origin to the target as $\theta\in[0,\frac{\pi}{2}]$ [^1] Note that we only consider $\theta\in[0,\frac{\pi}{2}]$ here. This is because in the case of $\theta\in(\frac{\pi}{2},\pi]$, we can redefine the origin as the ending position of the MA array, which yields the same sensing result thanks to symmetry., while the AoA is defined as $u=\cos\theta\in[0,1]$. Denote $\boldsymbol{s}_{n}=[x_{n},0]^{\top}$ as the coordinate of the $n$-th MA, and $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target, where $r_{\text{min}}$ and $r_{\text{max}}$ are the prescribed lower- and upper-bounds of the distance, respectively. Thus, the coordinate of the target is given by $\boldsymbol{r}=[r\cos\theta,r\sin\theta]^{\top}$. Then, the distance from the $n$-th MA to the target can be expressed as a function of the APV $\boldsymbol{x}$ and the target parameter vector denoted by $\boldsymbol{\eta}=[u,r]^{\top}$, i.e.,
@@ III MA-Enhanced Near-Field Sensing For 1D Arrays
$\hat{u}=\arg\max_{\bar{u}\in[0,1]}\frac{1}{\boldsymbol{\alpha}(\boldsymbol{x},\bar{u})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x},\bar{u})},$ (9)
@@
$\min_{\boldsymbol{x}}\max_{u\in[0,1]}\text{CRB}_{u}(\boldsymbol{x},u)\iff\max_{\boldsymbol{x}}\min_{u\in[0,1]}F_{u}(\boldsymbol{x},u).$ (14)
@@
The proof is the same as [ma2024movable, Appendix A] and thus omitted here for brevity. ∎
@@
The worst-case CRB on the AoA estimation, i.e., $\text{CRB}_{u}(\boldsymbol{x},0)$, decays with $A$ in the order of ${\cal O}(A^{-2})$.
Proof: The proof is the same as [ma2024movable, Appendix A] and thus omitted here for brevity.
Theorem 1 demonstrates that, to minimize the CRB of the AoA estimation MSE in the near-field, the optimal MA positions are the same as those for the AoA estimation in the far-field, as derived in [ma2024movable]. In particular, the MAs should be divided into two groups, as depicted in Fig. 2. The first group of MAs is placed at the leftmost end of the 1D MA array, while the other group is at the rightmost end. The performance equivalence between near-field and far-field scenarios stems from prior knowledge of the target distance. Additionally, it can be shown from Corollary 1 that the CRB on the AoA can be effectively decreased by increasing the length of the MA array, as this results in a larger array aperture, enabling the synthesis of sensing beams with higher angular resolution in the near-field region for a given distance.
@@
Theorem 2.
The optimal solution to (P1-2) is given by (17) presented in Theorem 1.
Proof:
See Appendix C. ∎
In addition, we provide the following corollary to characterize the scaling law of the worst-case CRB on the distance w.r.t. the array aperture $A$.
Corollary 2.
The worst-case CRB on the distance estimation, i.e., $\text{CRB}_{r}(\boldsymbol{x},r_{\text{max}})$, decreases with $A$ in the order of ${\cal O}(A^{-4})$.
Proof:
See Appendix D. ∎
It is noteworthy that the optimal APVs are identical in Cases 1.1 and 1.2 for estimating the AoA and distance, respectively. Both APVs maximally increase the aperture to ensure sensing resolution. In addition, it is noted from Corollary 2 that the worst-case CRB on the distance decays faster than that on the AoA ($A^{-4}$ versus $A^{-2}$). This implies that increasing the array aperture sharpens the near-field focal depth more significantly than it narrows the angular main lobe. The reason is that angle estimation depends on the linear phase gradient in (3) across the array, whereas distance estimation relies on the quadratic phase variation in (3) associated with wavefront curvature.
III-C Joint AoA and Distance Estimation in Case 1.3
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,1]\times[r_{\text{min}},r_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\boldsymbol{x},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x},\bar{\boldsymbol{\eta}})}.$ (22)
respectively. The procedures for deriving the CRBs are provided in Appendix E. It can be observed from (23) and (24) that there may exist a fundamental trade-off between minimizing $\text{CRB}_{u}(\boldsymbol{x})$ and $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ due to the complicated coupling between the variance and covariance terms therein. Moreover, although $\text{CRB}_{u}(\boldsymbol{x})$ is independent of the AoA, $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ depends on both the AoA and distance. To overcome this difficulty, we adopt the sum of $\text{CRB}_{u}(\boldsymbol{x})$ and the worst-case $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, i.e., $\text{CRB}_{u}(\boldsymbol{x})+\max_{\boldsymbol{\eta}}\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, as a performance metric to optimize the APV $\boldsymbol{x}$. As shown in (24), $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ is constantly positive and decreases monotonically with both $u$ and $r$. Thus, it reaches the maximum at the maximum values of $u$ and $r$, leading to the estimator that yields the worst-case CRB on the distance $\boldsymbol{\eta}_{\text{opt}}=[1,r_{\text{max}}]$. Then, the associated optimization problem can be equivalently transformed into the following problem, i.e.,
[…]
@@ IV-A System Model
Let the center of the square plane be the origin and denote the coordinate of the $n$-th MA by $\boldsymbol{s}_{n}\triangleq[x_{n},y_{n}]^{\top}\in{\cal G}$. Then, the coordinates of the MAs should satisfy $x_{n},y_{n}\in[-\frac{A}{2},\frac{A}{2}],n\in{\cal N}$. The APM of all $N$ MAs is denoted by $\tilde{\boldsymbol{s}}\triangleq[\boldsymbol{s}_{1},\boldsymbol{s}_{2},\ldots,\boldsymbol{s}_{N}]\in\mathbb{R}^{2\times N}$. As depicted in Fig. 4, the elevation and azimuth AoAs of the LoS path from the origin of the MA array to the target are denoted by $\theta\in[0,\frac{\pi}{2}]$ and $\phi\in[0,\frac{\pi}{2}]$ [^2] For analytical simplicity, we only consider the case where the target is in the first quadrant of the plane, i.e., $\theta\in[0,\frac{\pi}{2}]$ and $\phi\in[0,\frac{\pi}{2}]$. In the case that the target is in the other three quadrants, we can flip the MAs along the x/y-axis and apply our proposed algorithm accordingly., respectively. For analytical convenience, the two AoAs are respectively defined as
$u\triangleq\sin\theta\cos\phi\in[0,1],\quad v\triangleq\cos\theta\in[0,1].$ (28)
Then, the unit wave vector of the LoS path can be written as $\boldsymbol{n}=[u,v,\sqrt{1-u^{2}-v^{2}}]^{\top}$. Denote $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target. Hence, the positioning vector of the target is $\boldsymbol{r}=r\boldsymbol{n}$, and the distance from the $n$-th MA to the target can be expressed as a function of the APM $\tilde{\boldsymbol{s}}$ and the target parameters, i.e.,
@@
$\small h_{n}(\boldsymbol{s}_{n},\boldsymbol{\eta})=\beta\exp\Big(j\frac{2\pi}{\lambda}\big(x_{n}u+y_{n}v+\frac{(x_{n}^{2}+y_{n}^{2})(x_{n}u+y_{n}v)^{2}}{2r}\big)\Big).$ (30)
@@
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,1]\times[0,1]}\frac{1}{\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})},$ (32)
which can be obtained by performing an exhaustive search for $\bar{\boldsymbol{\eta}}=[\bar{u},\bar{v}]^{\top}$ over the interval $0,1]\times[0,1]$. Based on the 2D-MUSIC algorithm, the CRB matrix of the estimator vector can be derived. For notational simplicity, we define
@@
(a) $\theta=90^{\circ}$, $\phi=90^{\circ}$
(b) $\theta=0^{\circ}$, $\phi=0^{\circ}$
Figure 5: Target’s elevation and azimuth angles in the broadside and end-fire directions for the UPA with a maximum aperture.
@@
Figure 6: Target’s elevation and azimuth angles under the worst sensing case for the proposed MA array and $G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\boldsymbol{\eta})$ versus $\theta$ and $\phi$.
@@
$\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,1]\times[0,1]}G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}}),$ (37)
@@
To circumvent this difficulty, we first consider a uniform planar array (UPA), where the spacing between any two adjacent MAs is set as $A/(\sqrt{N}-1)$ to achieve the maximum aperture. We also assume a target positioned in the broadside direction of the UPA, i.e., both its elevation and azimuth angles are $90^{\circ}$, corresponding to $u=v=0$, as shown in Fig. 5(a). Notably, this leads to the minimum variation in the path length from each antenna element to the target under small perturbations in the horizontal or vertical direction. In contrast, if either the elevation or the azimuth angle approaches $0^{\circ}$, as shown in Fig. 5(b), even a slight angular variation can cause significant and uneven path length changes across the array aperture. This indicates that for any symmetric geometry (e.g., UPA), the worst-case sum of the CRBs on the AoAs occur at $u=v=0$, because the array response is least sensitive to angular changes in that geometry.
However, for an asymmetric array geometry, it is generally difficult to derive the two AoAs that yield the worst-case sum of the CRBs as for UPA. In fact, for the asymmetric array geometry shown in Fig. 6(a), it can be shown that the worst-case sum of the CRBs is achieved at $\theta=86.4^{\circ}$ and $\phi=90^{\circ}$ by employing a 2D exhaustive search. Nevertheless, as shown in Fig. 6(b), the difference in $G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})$ between its global maximum and the broadside direction remains below 0.1%. In addition to the geometry shown in Fig. 6(a), we have also checked other array geometries and consistently observed small performance differences. The details are omitted due to the page limit. Therefore, we put forward the following conjecture: for an MA array with the maximum aperture, the worst-case sensing performance approximately occurs when the target is located in the broadside direction, i.e., $\boldsymbol{\eta}_{\text{bd}}\triangleq[0,0]^{\top},\forall\tilde{\boldsymbol{s}}$. Importantly, this conjecture can greatly simplify the optimization problem while ensuring both mathematical tractability and minimal performance compromise.
In fact, as will be shown in Fig. 12(a) in Section V, the optimized array geometry for Case 2.1 is symmetric in both dimensions, which substantiates this conjecture. Consequently, we set $\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=\boldsymbol{\eta}_{\text{bd}}=[0,0]^{\top},\forall\tilde{\boldsymbol{s}}$, and the remaining MA position optimization problem becomes
@@
Specifically, we uniformly sample the continuous 2D MA array into $\tilde{M}\triangleq M^{2}$ $(\tilde{M}\gg N)$ discrete points, with $M$ points along each dimension. Hence, the inter-point spacing in each dimension is given by $\delta_{s}=\frac{A}{M}$, and the coordinate of the $(k,l)$-th sampling point is $\boldsymbol{s}_{kl}=[k\delta_{s},l\delta_{s}]^{\top},k,l\in{\cal M}\triangleq\{\frac{-M+1}{2},\frac{-M+3}{2},\ldots,\frac{M-3}{2},\frac{M-1}{2}\}$. The set of all sampling points is denoted as ${\cal Q}=\{\boldsymbol{s}_{kl}|k,l\in{\cal M}\}$. Based on this, we initialize the position set as ${\cal Q}_{\text{init}}=\{\boldsymbol{s}^{\text{init}}_{n}|\boldsymbol{s}^{\text{init}}_{n}\in{\cal Q},n\in{\cal N}\}$. In the $n$-th iteration, the algorithm exclusively adjusts the position of the $n$-th MA, i.e., $\boldsymbol{s}^{\text{init}}_{n}$, while fixing the positions of the remaining (N1)(N-1) MAs. Denote $\boldsymbol{s}^{\star}_{n}$ as the updated position of the $n$-th MA in the $n$-th iteration. Then, the set of all feasible sampling points for updating $\boldsymbol{s}^{\text{init}}_{n}$ is
$n+1\leq j\leq N\},2\leq n\leq N-1,$ (39)
@@
(P2-2) is equivalent to maximizing the variations among the $N$ terms in $\boldsymbol{\rho}$ which are correlated with the coordinates of the MAs. Due to the complex objective function, a closed-form optimal solution to (P2-2) cannot be derived similarly as in Case 1.2. As a result, we adopt the discrete-sampling based algorithm again to obtain a high-quality suboptimal solution. The procedures are similar to those of Algorithm 2 and thus omitted.
@@
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,1]\times[0,1]\times[r_{\text{min}},r_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})}.$ (44)
@@
Similar to Case 1.3, we aim to minimize the worst-case sum of the above three CRBs w.r.t. the estimators. On the one hand, based on the conjecture made in Case 2.1, $u=v=0$ yields the worst-case sum of the CRBs on the two AoAs estimation. On the other hand, $r_{\text{max}}$ yields the worst-case CRB on the distance estimation, as analyzed in Case 2.2. Consequently, the joint estimation performance reaches its worst-case condition when the target is located exactly above the origin of the plane with the maximum distance. The estimator vector that yields the worst-case sum of the CRBs is set as $\boldsymbol{\eta}_{\text{opt}}=[0,0,r_{\text{max}}]^{\top}$. The corresponding optimization problem can be expressed as
@@
$\quad+\widetilde{\text{CRB}}_{r}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{opt}})\Big)^{-1}$ (48a)

### v2

@@ II System Model
During the sensing process, the MA array transmits sensing signals and receives the echoes reflected from the target, which is assumed to remain static throughout the process[wang2023near, liu2022cramer]. To characterize the near-field channel from the antenna array to the target, we adopt the uniform spherical wave (USW) channel model in [liu2023near], where the channel coefficients have identical amplitudes across all MAs, while their phases vary across them. As depicted in Fig. 1, we denote the physical steering angle between the origin and the target as $\theta$, with $\theta\in[\theta_{\text{min}},\frac{\pi}{2}]$, where $\theta_{\text{min}}>0$ is the prescribed lower-bound of the steering angle. [^1] Note that we only consider $\theta\leq\frac{\pi}{2}$ here. This is because in the case of $\theta>\frac{\pi}{2}$, we can redefine the origin as the ending position of the MA array, which yields the same sensing result thanks to symmetry. Accordingly, the directional cosine of the AoA is defined as $u=\cos\theta\in[0,u_{\text{max}}]$, where $u_{\text{max}}=\cos\theta_{\text{min}}$ is its upper bound. Denote $\boldsymbol{s}_{n}=[x_{n},0]^{\top}$ as the coordinate of the $n$-th MA, and $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target, where $r_{\text{min}}$ and $r_{\text{max}}$ are the prescribed lower- and upper-bounds of the distance, respectively. Thus, the coordinate of the target is given by $\boldsymbol{r}=[r\cos\theta,r\sin\theta]^{\top}$. Then, the distance from the $n$-th MA to the target can be expressed as a function of the APV $\boldsymbol{x}$ and the target parameter vector denoted by $\boldsymbol{\eta}=[u,r]^{\top}$, i.e.,
@@ III MA-Enhanced Near-Field Sensing For 1D Arrays
$\hat{u}=\arg\max_{\bar{u}\in[0,u_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\boldsymbol{x},\bar{u})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x},\bar{u})},$ (9)
@@
$\min_{\boldsymbol{x}}\max_{u\in[0,u_{\text{max}}]}\text{CRB}_{u}(\boldsymbol{x},u)\iff\max_{\boldsymbol{x}}\min_{u\in[0,u_{\text{max}}]}F_{u}(\boldsymbol{x},u).$ (14)
@@
Note that problem (P1-1) has the same form as (P1) in [ma2024movable]. Therefore, its optimal APV can be derived by following the same procedures as those in [ma2024movable, Appendix A]. For brevity, the details are omitted. ∎
@@
The associated worst-case CRB on the AoA estimation for the optimal APV, i.e., $\text{CRB}_{u}(\boldsymbol{x}^{\star},0)$, decreases with $A$ in the order of ${\cal O}(A^{-2})$.
Proof:
Corollary 1 can be readily verified by substituting (17) into the worst-case CRB, i.e., $\text{CRB}_{u}(\boldsymbol{x},0)$. It follows that $\text{CRB}_{u}(\boldsymbol{x}^{\star},0)$ (given in [ma2024movable, (52)] and omitted here for brevity) decreases with $A$ for $A\geq(N-1)d$ in the order of ${\cal O}(A^{-2})$. This completes the proof. ∎
Theorem 1 demonstrates that, to minimize the CRB of the AoA estimation MSE in the near-field, the optimal MA positions are the same as those for the AoA estimation in the far-field, as derived in [ma2024movable]. In particular, the MAs should be divided into two groups, as depicted in Fig. 2. The first group of MAs is placed at the leftmost end of the 1D MA array, while the other group is at the rightmost end. The performance equivalence between near-field and far-field scenarios stems from prior knowledge of the target distance. Additionally, it can be shown from Corollary 1 that the CRB on the AoA estimation for the optimal APV can be effectively decreased by increasing the length of the MA array, as this results in a larger array aperture, enabling the synthesis of sensing beams with higher angular resolution in the near-field region for a given distance.
@@
Theorem 2.
The optimal solution to (P1-2) is given by (17) presented in Theorem 1.
Proof:
See Appendix C. ∎
In addition, we provide the following corollary to characterize the scaling law of the worst-case CRB on the distance estimation for the optimal APV, i.e., ${\text{CRB}}_{r}(\boldsymbol{x}^{\star},r_{\text{max}})$, w.r.t. the array aperture $A$.
Corollary 2.
The associated worst-case CRB on the distance estimation for the optimal APV, i.e., $\text{CRB}_{r}(\boldsymbol{x}^{\star},r_{\text{max}})$, decreases with $A$ in the order of ${\cal O}(A^{-4})$.
Proof:
See Appendix D. ∎
It is noteworthy that the optimal APVs are identical in Cases 1.1 and 1.2 for estimating the AoA and distance, respectively. Both APVs maximally increase the aperture to ensure sensing resolution. In addition, it is noted from Corollary 2 that the worst-case CRB on the distance estimation for the optimal APV decays faster than that on the AoA ($A^{-4}$ versus $A^{-2}$). This implies that increasing the array aperture sharpens the near-field focal depth more significantly than it narrows the angular main lobe. The reason is that angle estimation depends on the linear phase gradient in (3) across the array, whereas distance estimation relies on the quadratic phase variation in (3) associated with wavefront curvature.
III-C Joint AoA and Distance Estimation in Case 1.3
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,u_{\text{max}}]\times[r_{\text{min}},r_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\boldsymbol{x},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x},\bar{\boldsymbol{\eta}})}.$ (22)
respectively. The procedures for deriving the CRBs are provided in Appendix E. It can be observed from (23) and (24) that there may exist a fundamental trade-off between minimizing $\text{CRB}_{u}(\boldsymbol{x})$ and $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ due to the complicated coupling between the variance and covariance terms therein. Moreover, although $\text{CRB}_{u}(\boldsymbol{x})$ is independent of the AoA, $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ depends on both the AoA and distance. To overcome this difficulty, we adopt the sum of $\text{CRB}_{u}(\boldsymbol{x})$ and the worst-case $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, i.e., $\text{CRB}_{u}(\boldsymbol{x})+\max_{\boldsymbol{\eta}}\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, as a performance metric to optimize the APV $\boldsymbol{x}$. It is noted from (24) that $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ is constantly positive and increases monotonically with both $u$ and $r$. Thus, it reaches the maximum at the maximum values of $u$ and $r$. As such, the estimator vector that yields the worst-case $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ is given by $\boldsymbol{\eta}_{\text{opt}}=[u_{\text{max}},r_{\text{max}}]^{\top}$. This indicates that, in Case 1.3, the worst-case performance arises when the target lies closest to the end-fire direction and farthest from the MA array. Hence, the associated optimization problem can be formulated as
[…]
@@ IV-A System Model
Let the center of the square plane be the origin and denote the coordinate of the $n$-th MA by $\boldsymbol{s}_{n}\triangleq[x_{n},y_{n}]^{\top}\in{\cal G}$. Then, the coordinates of the MAs should satisfy $x_{n},y_{n}\in[-\frac{A}{2},\frac{A}{2}],n\in{\cal N}$. The APM of all $N$ MAs is denoted by $\tilde{\boldsymbol{s}}\triangleq[\boldsymbol{s}_{1},\boldsymbol{s}_{2},\ldots,\boldsymbol{s}_{N}]\in\mathbb{R}^{2\times N}$. As depicted in Fig. 4, the elevation and azimuth steering angles of the LoS path from the origin of the MA array to the target are denoted by $\theta\in[\theta_{\text{min}},\frac{\pi}{2}]$ and $\phi\in[\phi_{\text{min}},\frac{\pi}{2}]$, respectively, where $\theta_{\text{min}}>0$ and $\phi_{\text{min}}>0$ are the prescribed lower-bounds of the elevation and azimuth AoAs, respectively. [^2] For analytical simplicity, we only consider the case where the target is in the first quadrant of the plane, i.e., $\theta\in(0,\frac{\pi}{2}]$ and $\phi\in(0,\frac{\pi}{2}]$. In the case that the target is in the other three quadrants, we can flip the MAs along the xx/$y$-axis and apply our proposed algorithm accordingly. For analytical convenience, the two AoAs are respectively defined as
$u\triangleq\sin\theta\cos\phi\in[0,u_{\text{max}}],\quad v\triangleq\cos\theta\in[0,v_{\text{max}}],$ (28)
where $u_{\text{max}}$ and $v_{\text{max}}$ are their upper-bounds, respectively. Then, the unit wave vector of the LoS path can be written as $\boldsymbol{n}=[u,v,\sqrt{1-u^{2}-v^{2}}]^{\top}$. Denote $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target. Hence, the positioning vector of the target is $\boldsymbol{r}=r\boldsymbol{n}$, and the distance from the $n$-th MA to the target can be expressed as a function of the APM $\tilde{\boldsymbol{s}}$ and the target parameters, i.e.,
@@
$h_{n}$ $(\boldsymbol{s}_{n},\boldsymbol{\eta})=$ (30)
$\beta\exp\Big(j\frac{2\pi}{\lambda}\big(x_{n}u+y_{n}v+\frac{(x_{n}^{2}+y_{n}^{2})-(x_{n}u+y_{n}v)^{2}}{2r}\big)\Big).$
@@
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,u_{\text{max}}]\times[0,v_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})},$ (32)
which can be obtained by performing an exhaustive search for $\bar{\boldsymbol{\eta}}=[\bar{u},\bar{v}]^{\top}$ over the interval $0,u_{\text{max}}]\times[0,v_{\text{max}}]$. Based on the 2D-MUSIC algorithm, the CRB matrix of the estimator vector can be derived. For notational simplicity, we define
@@
Figure 5: Target’s elevation and azimuth angles under the worst sensing case for an asymmetric APM.
@@
$\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,u_{\text{max}}]\times[0,v_{\text{max}}]}G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}}),$ (37)
@@
To circumvent this difficulty, we first consider a symmetric array geometry or APM. Inspired by Case 1.1, we expect that the worst-case performance for AoA estimation for the 2D array also occurs when the target is in the broadside direction, i.e., both its elevation and azimuth angles are $90^{\circ}$, corresponding to $u=v=0$. Rigorously, for any symmetric APM $\tilde{\boldsymbol{s}}$, the results in (35) and (36) indicate that $G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\boldsymbol{\eta})$ achieves the maximum when both $\text{var}(\boldsymbol{\xi})$ and $\text{var}(\boldsymbol{\pi})$ reach the minimum while $\text{cov}(\boldsymbol{\xi},\boldsymbol{\pi})$ reaches the maximum. From (33) and (34), it is noted that the variations among the $N$ terms in $\boldsymbol{\xi}$ and $\boldsymbol{\pi}$ are minimal when $u=v=0$, which yields the minimum $\text{var}(\boldsymbol{\xi})$ and $\text{var}(\boldsymbol{\pi})$. Moreover, when $u=v=0$, $\boldsymbol{\xi}$ and $\boldsymbol{\pi}$ degrade to $\boldsymbol{x}=[x_{1},x_{2},\ldots,x_{N}]^{\top}$ and $\boldsymbol{y}=[y_{1},y_{2},\ldots,y_{N}]^{\top}$, respectively. Since $\boldsymbol{x}$ is identical to $\boldsymbol{y}$ for any symmetric APM, the linear correlation between the $N$ terms in $\boldsymbol{\xi}$ and $\boldsymbol{\pi}$ is maximized, which in turn yields the maximum $\text{cov}(\boldsymbol{\xi},\boldsymbol{\pi})$. It follows from the above that $\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=[0,0]^{\top}$ holds for symmetric array geometries.
However, for an asymmetric array geometry or APM, it is generally difficult to derive the two AoAs that yield the worst-case sum of the CRBs. In fact, for the asymmetric array geometry shown in Fig. 5(a), it can be shown that the worst-case sum of the CRBs is achieved at $\theta=86.4^{\circ}$ and $\phi=90^{\circ}$ by employing a 2D exhaustive search. Nevertheless, as shown in Fig. 5(b), the difference in $G_{\boldsymbol{\eta}}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})$ between its global maximum and the broadside direction remains below 1%. In addition to the geometry shown in Fig. 5(a), we have also checked other array geometries and consistently observed small performance differences. The details are omitted due to the page limit. Therefore, we set $\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=\boldsymbol{\eta}_{\text{bd}}\triangleq[0,0]^{\top},\forall\tilde{\boldsymbol{s}}$, which greatly simplifies the optimization problem while ensuring both mathematical tractability and minimal performance compromise. In fact, as will be shown in Fig. 11(a) in Section V via simulation, the optimized array geometry for Case 2.1 is symmetric in both dimensions.
With $\boldsymbol{\eta}_{\text{opt}}(\tilde{\boldsymbol{s}})=[0,0]^{\top},\forall\tilde{\boldsymbol{s}}$, the remaining MA position optimization problem becomes
@@
Specifically, we uniformly sample the continuous 2D MA array into $\tilde{M}\triangleq M^{2}$ $(\tilde{M}\gg N)$ discrete points, with $M$ points along each dimension. Hence, the inter-point spacing in each dimension is given by $\delta_{s}=\frac{A}{M}$, and the coordinate of the $(k,l)$-th sampling point is $\boldsymbol{s}_{kl}=[k\delta_{s},l\delta_{s}]^{\top},k,l\in\bar{{\cal M}}\triangleq\{\frac{-M+1}{2},\frac{-M+3}{2},\ldots,\frac{M-3}{2},\frac{M-1}{2}\}$. The set of all sampling points is denoted as ${\cal Q}=\{\boldsymbol{s}_{kl}|k,l\in\bar{{\cal M}}$. Based on this, we initialize the position set as ${\cal Q}_{\text{init}}=\{\boldsymbol{s}^{\text{init}}_{n}|\boldsymbol{s}^{\text{init}}_{n}\in{\cal Q},n\in{\cal N}\}$. In the $n$-th iteration, the algorithm exclusively adjusts the position of the $n$-th MA, i.e., $\boldsymbol{s}^{\text{init}}_{n}$, while fixing the positions of the remaining (N1)(N-1) MAs. Denote $\boldsymbol{s}^{\star}_{n}$ as the updated position of the $n$-th MA in the $n$-th iteration. Then, the set of all feasible sampling points for updating $\boldsymbol{s}^{\text{init}}_{n}$ is
$n+1\leq j\leq N\},2\leq n\leq N-1,$ (39)
@@
(P2-2) is equivalent to maximizing the variations among the $N$ terms in $\boldsymbol{\rho}$ that are correlated with the coordinates of the MAs. Due to the complex objective function, a closed-form optimal solution to (P2-2) cannot be derived similarly as in Case 1.2. As a result, we adopt the discrete-sampling based algorithm again to obtain a high-quality suboptimal solution. The procedures are similar to those of Algorithm 2 and thus omitted.
@@
$\hat{\boldsymbol{\eta}}=\arg\max_{\bar{\boldsymbol{\eta}}\in[0,u_{\text{max}}]\times[0,v_{\text{max}}]\times[r_{\text{min}},r_{\text{max}}]}\frac{1}{\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})^{\mathsf{H}}\boldsymbol{U}_{\boldsymbol{w}}\boldsymbol{U}_{\boldsymbol{w}}^{\mathsf{H}}\boldsymbol{\alpha}(\tilde{\boldsymbol{s}},\bar{\boldsymbol{\eta}})}.$ (44)